OpenStudy (anonymous):

theta where 0 degrees le theta < 360 degrees , to the nearest tenth of a degree

OpenStudy (mathmale):

Sorry, but something seems missing from your problem statement. Exactly what are you looking for?

OpenStudy (anonymous):

\[1. \cos \theta = 0.5563 \]

OpenStudy (jdoe0001):

hmm the who? that reads like \(\bf \theta=?\qquad 0\le \theta< 360^o\) which hmm leaves us in the lurch

OpenStudy (jdoe0001):

?

OpenStudy (anonymous):

Sorry let me rewrite the question

OpenStudy (jdoe0001):

well, you can always just take the inverse function of it and get the angle out \(\bf csc(\theta)=3.2745\implies csc^{-1}[csc(\theta)]=csc^{-1}(3.2745) \\ \quad \\ \theta=csc^{-1}(3.2745)\)

OpenStudy (jdoe0001):

hmmmm lemme see if I can rephrase that some

OpenStudy (jdoe0001):

\(\bf csc(\theta)=3.2745\implies \cfrac{1}{sin(\theta)}=3.2745 \\ \quad \\ \cfrac{1}{3.2745}=sin(\theta)\implies sin^{-1}\left( \cfrac{1}{3.2745} \right)=sin^{-1}[sin(\theta)] \\ \quad \\ sin^{-1}\left( \cfrac{1}{3.2745} \right)=\theta\)

OpenStudy (jdoe0001):

hmm it'd be easier if you post a screenshot of the material so we can see what's being asked :)

OpenStudy (anonymous):

i'll try that :)

OpenStudy (mathmale):

Please post just one question at a time. Having to read through an entire conversation looking for the problem at hand consumes precious time. Thank you.